OpenStudy (anonymous):

The base of a hemisphere has an area of 256π cm2. Find its volume.

OpenStudy (anonymous):

pi*r^2 = 256*pi (area of the circle which is the base of a hemisphere) so r = 16 Now volume of a hemisphere = (4/3*pi*r^3)/2.Put r= 16 in this and you will get your volume